Ethylene glycol (C2H6O2) is a non-volatile and non-electrolyte compound with a density of 1.11 g.cm–3 at 25.0oC. Calculate the volume ethylene glycol that must be added to 675 cm3 of water to reduce the vapour pressure by 3.82% at 25.0oC. The density and vapour pressure of pure water at 25.0oC are 1.02 g.cm–3 and  

23.8 torr respectively



1
Expert's answer
2021-03-29T07:28:45-0400

Q176319

Ethylene glycol (C2H6O2) is a non-volatile and non-electrolyte compound with a density of 1.11 g.cm–3 at 25.0oC. Calculate the volume ethylene glycol that must be added to 675 cm3 of water to reduce the vapor pressure by 3.82% at 25.0oC. The density and vapor pressure of pure water at 25.0oC are 1.02 g.cm–3 and  

23.8 torr respectively.


Solution:


Step 1 : To find the reduced vapor pressure of water.


Vapor pressure of water at 25.0 0C = 23.8 torr.


We have to find the volume of ethylene glycol that must be added to 675 cm3 of water, so as to reduce the vapor pressure by 3.82 % .


New vapor pressure will be equal = 100 - 3.82 = 96.18 % of initial vapor pressure.


96.18 % of 23.8 torr


"= 23.8 torr * \frac{96.18}{100} = 22.891 torr."



Hence the reduced vapor pressure of water would be 22.89 torr.


Step 2 : To find the mole fraction of water when its vapor pressure is reduced to 22.89 torr.


Let P1 be the vapor pressure of water in the solution and

P10 be the vapor pressure of pure water.


Then the vapor pressure of water in the solution is given by the expression


"P_1 = x_1 * P_1 ^0 ."


where x1 is the mole fraction of water.


substitute P1 = 22.89 torr and P10 = 23.8 torr, in the formula we have


"22.89\space torr = x_1 * 23.8\space torr."


divide both the side by 23.8 torr, we have


"x_1 = \frac{22.89\space torr }{23.8\space torr} = 0.9618 ."



Step 3 : To find the moles of water in the solution.


volume of water = 675 cm3 .

density of water = 1.02 g/cm3 .


plug volume and density of water in the density formula and find the mass of the water.


"density = \frac{mass}{volume } ;"


arranging this formula for mass we have


"mass = density * volume = 1.02\space g/cm^3 * 675\space cm^3 = 688.5\space grams."


molar mass of water = 18.0148 g/mol.




"moles\space of\space water = 688.5\space g\space of\space H_2O * \frac{1\space mol\space of \space H_2O}{18.0148\space g\space of\space H_2O}"

"= 38.219\space mol\space of \space H_2O"


Step 4 : To find the moles of moles of ethylene glycol.



"mole\space fraction\space of\space water = \frac{mol\space of \space water }{moles\space of\space solution }"

"total \space moles \space in \space solution = \frac{mol \space of \space water }{mole \space fraction \space of \space water }"


"total \space moles \space in \space solution = \frac{38.219\space mol }{0.9618 } = 39.737\space moles"



moles of ethylene glycol = moles of solution - moles of solvent

= 39.737 mol - 38.219 mol

= 1.5180 moles of ethylene glycol.


Step 5 : To convert 1.518 mol of ethylene glycol to grams.


molar mass of ethylene glycol = 62.0674 g/mole


grams of ethylene glycol = moles * molar mass

= 1.5180 mol * 62.0674 g/mol

= 94.2121 grams.


Step 6 : To convert 94.2121 grams of ethylene glycol to 'mL'



volume = mass/ density = 94.2121/ 1.11 g/mL

= 84.89 mL


Which in 3 significant figure is 84.9 mL


Hence we must add 84.9 mL of ethylene glycol to 675 mL of water, so as to reduce the vapor pressure by 3.82 % of initial vapor pressure of water.